OpenStudy (anonymous):

HELP!!! The surface area of two similar rectangular prisms are 361m2 and 441m2. What is the scale factor of the larger prism to the smaller prism? A.) 361/441 B.) 19/21 C.) 21/19 D.) 19/11

OpenStudy (anonymous):

@ganeshie8 please help

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@nincompoop

ganeshie8 (ganeshie8):

Hint : If areas of similar figures are in ratio \(\dfrac{a}{b}\), then the lengths(scale factors) will be in ratio \(\dfrac{\sqrt{a}}{\sqrt{b}}\)

OpenStudy (anonymous):

so i just take each answer choice and do that to them??

OpenStudy (anonymous):

i'm still not understanding

ganeshie8 (ganeshie8):

Yes, first setup the ratio of areas

OpenStudy (anonymous):

is it 19/21

ganeshie8 (ganeshie8):

careful, they want larger prism on top

ganeshie8 (ganeshie8):

ratio of areas = \(\dfrac{441}{361}\)

OpenStudy (anonymous):

so its C?

ganeshie8 (ganeshie8):

so, ratio of scalefactors = \(\dfrac{\sqrt{441}}{\sqrt{361}}=?\)

OpenStudy (anonymous):

21/19

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

i have another question if thats ok

ganeshie8 (ganeshie8):

sure ask

OpenStudy (anonymous):

The two cones are similar. Find the surface area of the larger cone. Round your answer to the nearest hundredth. A. 810.28 B. 847.80 C. 868.30 D. 886.25

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

S=π×11.12+π×11.1×13.8 is the formula, solve that and u will get the answer

OpenStudy (anonymous):

I got 516..../

OpenStudy (anonymous):

that's not right

OpenStudy (anonymous):

hmm let me solve and see what i get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i am sorry the same question is on open study and I checked the formula I am so sorry I was not able to help

OpenStudy (anonymous):

i feel bad :(

OpenStudy (anonymous):

its ok. @ganeshie8 can you help

OpenStudy (anonymous):

thnx